# NewReff.bib

### v1

@@ 3 A Borel field of pairwise non-isomorphic groups
Let $\mathcal{Y}=\mathbb{N}^{\mathbb{N}}$ be the Baire space with the product topology. It is well known that $\mathcal{Y}$ is a Polish space and hence has a standard Borel structure. It is also well known that there is a continuous order-preserving bijection from $\mathcal{Y}$ to $(0,\infty)$. Hence, with respect to the Borel $\sigma$-algebras on both sides, this bijection is Borel. We shall usually denote sequences in $\mathbb{N}^{\mathbb{N}}$ by $U=(u_{n})_{n\in\mathbb{N}}$. For positive integers $k,l$, consider the subset $\mathcal{Y}_{k,l}\subset\mathcal{Y}$ given by $\{U\in\mathcal{Y}\;|\;u_{k}=l\}$, i.e., all sequences whose $k^{\text{th}}$ entry is $l$. Also consider the subset $\mathcal{Y}_{0}\subset\mathcal{Y}$ consisting of all strictly increasing sequences of odd integers $\geq 5$.
@@
For a positive integer $i$, let $\pi_{i}:\mathcal{Y}\rightarrow\mathbb{N}$ be the map $U\mapsto u_{i}$. By definition the product topology on $\mathcal{Y}$ is the weakest topology for which $\pi_{i}$ is continuous for all $i$. Since $\mathcal{Y}_{k,l}=\pi_{k}^{-1}(\{l\})$, it is an open set and hence clearly Borel. Now let $\mathbb{N}_{even}$ be the set of even positive integers. Clearly $E=\bigcup_{i\in\mathbb{N}}\pi_{i}^{-1}({\mathbb{N}_{even}})$ is a countable union of open sets and is open and Borel. Then $E^{C}=\mathcal{Y}\backslash E$ is also Borel (note that $E^{C}$ consists of all sequences with only odd integers). For all pairs $i,j\in\mathbb{N}$ with $i\neq j$, we have that $(\pi_{i},\pi_{j}):E^{C}\rightarrow\mathbb{N}\times\mathbb{N}$ is continuous and hence Borel. Let $F_{i}$ be the subset of $E^{C}$ consisting of sequences $(u_{n})_{n\in\mathbb{N}}$ such that $u_{i}\geq u_{i+1}$. Then $F_{i}$ is Borel because it is the inverse image under $(\pi_{i},\pi_{i+1})$ of the Borel set $\{(m,n)\;|\;m\geq n\}$ in $\mathbb{N}\times\mathbb{N}$. Then we notice that $E^{C}\backslash\bigcup_{i\geq 1}F_{i}$ consisting of all strictly increasing sequences of odd integers is a Borel set. Now we can subtract the Borel sets $\pi_{1}^{-1}(\{3\})$ and $\pi_{1}^{-1}(\{1\})$ to get that $\mathcal{Y}_{0}$ is indeed a Borel set. ∎
@@
The sections $\theta_{i,j}$ of course have the property that for all $U\in Y$, the set $\{\theta_{i,j}(U)\;|\;i,j\in\mathbb{N}\}$ equals $X_{U}$. Also for two tuples $(i,j)\neq(k,l)$ of positive integers, one can check that the maps $Y\rightarrow[0,\infty)$ given by $U\mapsto d_{u}(\theta_{i,j}(U),\theta_{k,l}(U))$ where $d_{U}$ is the metric that gives the discrete topology, are all Borel, essentially because the maps $\pi_{i}$ are Borel. Then there is a unique standard Borel structure on $X$ that makes the map $\pi:X\rightarrow Y$ and the countable family of sections $\theta_{i,j}$ Borel, for example using [WoutersVaes24, Proposition 4.2].
@@
[…]
For two Borel fields of separable structures $H=(H_{z})_{z\in Z}$ and $K=(K_{z})_{z\in Z}$ and the canonical projections $\pi_{H}:H\rightarrow Z$ and $\pi_{K}:K\rightarrow Z$, we denote by $H\times_{\pi}K$ the Borel set $\{(h,k)\in H\times K\;|\;\pi_{H}(h)=\pi_{K}(k)\}$.
@@
We claim that $\phi(\mathfrak{G})$ is a Borel subset of $Y\times X^{\mathbb{N}\times\mathbb{N}}$. To see this, let $\mathcal{S}$ be the set of finite words generated by the elements $\alpha$, $\beta$ and their inverses. By Lemma 3.3, and the fact that composition and inverses of Borel bijections are Borel, we have that if $g\in\mathcal{S}$, then $g:X\rightarrow X$ is Borel. Now consider the Borel map $B_{g}:Y\rightarrow X^{\mathbb{N}\times\mathbb{N}}$ given by $U\mapsto g\cdot(\theta_{i,j}(U))_{i,j}$. For each $g\in\mathcal{S}$, since $B_{g}$ is a Borel map, the graph of $B_{g}$ is Borel (for example see [TakesakiVolume1, Lemma A.9]). Now since $\mathcal{S}$ is countable, we have that the union of the graphs of $\{B_{g}\;|\;g\in\mathcal{S}\}$ is a Borel subset of $Y\times X^{\mathbb{N}\times\mathbb{N}}$, which is precisely equal to $\phi(\mathfrak{G})$. Now by [WoutersVaes24, Lemma 4.1], there is a unique standard Borel structure on $\mathfrak{G}$ making the maps $\pi:\mathfrak{G}\rightarrow Y$ and the map $C:\mathfrak{G}\times_{\pi}X\rightarrow X$ Borel. By [WoutersVaes24, Proposition 4.2], we have that every element of $\mathcal{S}$ is a Borel section $Y\rightarrow\mathfrak{G}$. It follows that the multiplication, inverse and identity maps are Borel and that $\mathfrak{G}=(\mathfrak{G}_{U})_{U\in Y}$ is a Borel field of countable discrete groups. ∎
@@ 4 Construction of genuine ergodic groupoids
Recall the notion of a basis of a discrete measured groupoid from [BCDK24, Definition 3.2]. Roughly speaking, the idea is that a discrete measured groupoid admits a countable disjoint collection of non-null Borel bisections such that the union is the entire groupoid up to measure zero. It turns out that for ergodic groupoids, one can construct a basis consisting of global bisections. This can be seen for example by [Kec95, Exercise 18.15] and we state this in our context as the next proposition. Notice that for non ergodic groupoids this cannot be true. For example, if we have a disjoint union of two measured fields of groups consisting of finite and infinite groups almost everywhere respectively, then such a discrete measured groupoid can never admit a basis of global bisections.
@@
We first prove that 2 implies 1 by showing that if $\mu$ is atomic, then any discrete measured groupoid with associated equivalence relation $\mathcal{R}$ is a transformation groupoid. Since $\mathcal{R}$ is ergodic and $(X,\mu)$ is atomic, without loss of generality $\mathcal{R}$ is the full equivalence relation on an at most countable set $X$ and $\mu$ can be chosen to be the counting measure without loss of generality. Consequently, the isotropy groups $\Gamma_{x}$ are all isomorphic and let us denote this group by $\Gamma$.
Since $\mathcal{G}$ is amenable, by [Popa-Shlyakhtenko-Vaes20, Proposition 6.5], $\mathcal{G}$ can be written as a semi-direct product $\mathcal{R}\ltimes_{\delta}(\Gamma_{x})_{x\in X}$ where $\delta$ is an action of $\mathcal{R}$ on the measured field of groups $(\Gamma_{x})_{x\in X}$. Let $G$ be a group with the same cardinality as $X$ and suppose for convenience that the unit space of $\mathcal{G}$ is $G$ with the counting measure. Consider the action of $G\times\Gamma$ on $G$ given by $(h,\gamma)\cdot g=hg$. Consider the transformation groupoid $(G\times\Gamma)\ltimes G$ and the map $\phi:(G\times\Gamma)\ltimes G\rightarrow\mathcal{G}$ given by:
@@
We can now check that that this is groupoid homomorphism. First note that $(hg,(k,\lambda))(g,(h,\gamma))=(g,(kh,\lambda\gamma))$ and $\phi(g,(kh,\lambda\gamma))=((khg,g),\delta_{(g,e)}(\lambda\gamma))$. One can check now that:
$\phi(hg,(k,\lambda))\phi(g,(h,\gamma))$ $=((khg,hg)\delta_{(hg,e)}(\lambda))((hg,g),\delta_{(g,e)}(\gamma))$
@@
It is easy to note now that $\phi$ is a bijective and since multiplication and the action $\delta$ is Borel, $\phi$ is a Borel groupoid isomorphism.
Now we prove that 1 implies 2, essentially using Proposition 4.1 and Corollary 3.4. Suppose that $\mathcal{R}$ is an ergodic equivalence relation on the diffuse standard probability space $(X,\mu)$ and as in proposition 4.1 consider a countable family of bijective Borel functions $\theta_{n}:X\rightarrow X$ such that the global bisections $B_{n}=\{(\theta_{n}(x),x)\;|\;x\in X\}$ covers all of $\mathcal{R}$. Let $(G^{0}_{x})_{x\in X}$ be a measured field of finitely generated pairwise non-isomorphic groups (for example as in Corollary 3.4). Let $K$ be an infinite group and note that by [AnaPopa10, Proposition 1.3.10] the wreath product $G_{x}=G^{0}_{x}\wr K$ is an icc group for all $x\in X$. By Lemma 2.3, the family of groups $(G_{x})_{x\in X}$ is a measured field of discrete countable groups.
@@
To show that 2 implies 1, one simply notes that if $\mathcal{G}$ is a discrete measured groupoid of type I, then almost every ergodic groupoid in the ergodic decomposition of $\mathcal{G}$ is of type I, and hence the statement follows from Theorem 4.2. To show that 1 implies 2, without loss of generality we can assume that $\mathcal{R}=\mathcal{R}_{\text{ns}}$ as in Notation 2.4. Suppose that $(Z,\eta)$ is a standard probability space and $(\mathcal{R}_{z})_{z\in Z}$ is the ergodic decomposition of $\mathcal{R}$, where $\mathcal{R}_{z}$ is an ergodic equivalence relation on a standard probability space $(Z_{0},\nu_{z})$ for a.e. $z\in Z$. If $(Z,\eta)$ is atomic, then $\mathcal{R}$ is a disjoint union of countably many ergodic equivalence relations. It follows once again from Theorem 4.2 that there exists genuine groupoids with associated equivalence relation $\mathcal{R}$.
Now suppose that $(Z,\eta)$ is diffuse and let $\nu$ be the integral of the $(\nu_{z})_{z\in Z}$ with respect to $\eta$. We know that $\mathcal{R}$ can be written as an equivalence relation on $(Z_{0}\times Z,\nu)$ given by $(w,z)\sim(w^{\prime},z^{\prime})$ if and only if $z=z^{\prime}$ and $(w,w^{\prime})\in\mathcal{R}_{z}$. Consider a measured field $G=(G_{w})_{w\in Z_{0}}$ of discrete countable finitely generated groups such that they are pairwise non-isomorphic almost everywhere, as in Corollary 3.4. Consider the measured field of groups $H=(H_{(w,z)})_{(w,z)\in Z_{0}\times Z}$ defined as $H_{(w,z)}\coloneqq G_{w}$ for all $z\in Z$ and $w\in Z_{0}$. Since $\mathcal{R}$ is not necessarily ergodic, we dont necessarily have a basis of global bisections as in Proposition 4.1. However we can work around this issue as follows.

### v2

@@ 3 A Borel field of pairwise non-isomorphic groups
Let $\mathcal{Y}=\mathbb{N}^{\mathbb{N}}$ be the Baire space with the product topology. It is well known that $\mathcal{Y}$ is a Polish space and hence has a standard Borel structure. It is also well known that there is a continuous order-preserving bijection from $\mathcal{Y}$ to $(0,\infty)$. Hence, with respect to the Borel $\sigma$-algebras on both sides, this bijection is Borel. We shall denote sequences in $\mathbb{N}^{\mathbb{N}}$ by $U=(u_{n})_{n\in\mathbb{N}}$. For positive integers $k,l$, consider the subset $\mathcal{Y}_{k,l}\subset\mathcal{Y}$ given by $\{U\in\mathcal{Y}\;|\;u_{k}=l\}$, i.e., all sequences whose $k^{\text{th}}$ entry is $l$. Also consider the subset $\mathcal{Y}_{0}\subset\mathcal{Y}$ consisting of all strictly increasing sequences of odd integers $\geq 5$.
@@
For a positive integer $i$, let $\pi_{i}:\mathcal{Y}\rightarrow\mathbb{N}$ be the map $U\mapsto u_{i}$. By definition the product topology on $\mathcal{Y}$ is the weakest topology for which $\pi_{i}$ is continuous for all $i$. Since $\mathcal{Y}_{k,l}=\pi_{k}^{-1}(\{l\})$, it is an open set and hence clearly Borel. Now let $\mathbb{N}_{even}$ be the set of even positive integers. Clearly $E=\bigcup_{i\in\mathbb{N}}\pi_{i}^{-1}({\mathbb{N}_{even}})$ is a countable union of open sets and is open and Borel. Then $\EE^{C}=\mathcal{Y}\backslash E$ is also Borel (note that $E^{C}$ consists of all sequences with only odd integers). For all pairs $i,j\in\mathbb{N}$ with $i\neq j$, we have that $(\pi_{i},\pi_{j}):E^{C}\rightarrow\mathbb{N}\times\mathbb{N}$ is continuous and hence Borel. Let $F_{i}$ be the subset of $E^{C}$ consisting of sequences $(u_{n})_{n\in\mathbb{N}}$ such that $u_{i}\geq u_{i+1}$. Then $F_{i}$ is Borel because it is the inverse image under $(\pi_{i},\pi_{i+1})$ of the Borel set $\{(m,n)\;|\;m\geq n\}$ in $\mathbb{N}\times\mathbb{N}$. Then we notice that $E^{C}\backslash\bigcup_{i\geq 1}F_{i}$ consisting of all strictly increasing sequences of odd integers is a Borel set. Now we can subtract the Borel sets $\pi_{1}^{-1}(\{3\})$ and $\pi_{1}^{-1}(\{1\})$ to get that $\mathcal{Y}_{0}$ is indeed a Borel set. ∎
@@
One can check that the sections $\theta_{i,j}$ have the property that for all $U\in Y$, the set $\{\theta_{i,j}(U)\;|\;i,j\in\mathbb{N}\}$ equals $X_{U}$. Also for two tuples $(i,j)\neq(k,l)$ of positive integers, one can check that the maps $Y\rightarrow[0,\infty)$ given by $U\mapsto d_{u}(\theta_{i,j}(U),\theta_{k,l}(U))$ where $d_{u}$ is the metric that gives the discrete topology, are all Borel, essentially because the maps $\pi_{i}$ are Borel. Then there is a unique standard Borel structure on $X$ that makes the map $\pi:X\rightarrow Y$ and the countable family of sections $\theta_{i,j}$ Borel, for example using [VW25, Proposition 4.2].
@@
For $U\in Y$, let $X^{i}_{U}=\{x_{i,1},x_{i,2},...,x_{i,u_{i}}\}$ and let $X^{i}=\bigsqcup_{U\in Y}X^{i}_{U}$. We have that $X^{i}$ is a Borel subset of $X$ for all $i$ because $X^{i}=\bigcup_{j}\theta_{i,j}(Y)$. Now denote by $X^{i,j}$ the Borel subset $\pi^{-1}(Y_{i,j})$ of $X^{i}$. Notice that it is enough to prove that $\alpha|_{X^{i,j}}:X^{i,j}\rightarrow X^{i,j}$ is Borel for all $i,j\in\mathbb{N}$. We know that $\{\theta_{i,k}\;|\;1\leq k\leq j\}$ is a total set of Borel sections on $Y_{i,j}$. To check that $\alpha|_{X^{i,j}}$ is Borel, it is enough to prove that $\alpha\circ\theta_{i,k}$ is a Borel section for all $1\leq k\leq j$ by [Wou23, Lemma A.2.6]. But this holds because restricted to $X^{i,j}$, one has that $\alpha\circ\theta_{i,k}=\theta_{i,k+1}$ if $k\neq j$ and $\alpha\circ\theta_{i,j}=\theta_{i,1}$. Since $\{X^{i,j}\;|\;i,j\in\mathbb{N}\}$ forms a countable Borel partition of $X$, we have that $\alpha:X\rightarrow X$ is a countable disjoint union of Borel maps and is hence Borel.
Similarly one can consider the subsets $W^{i}_{U}=\{x_{i,1},x_{i,2},x_{i,3}\}$ and let $W^{i}$ be the disjoint union $\bigsqcup_{Y\in U}W^{i}_{U}$. Exactly as in the previous paragraph, $W^{i}$ is a Borel subset of $X$ for all $i\in\mathbb{N}$. One has that restricted to $W^{i}$, the set $\{\theta_{i,j}\;|\;1\leq j\leq 3\}$ is a total set of Borel sections, and clearly $\beta\circ\theta_{i,j}$ is again Borel. Once again by [Wou23, Lemma A.2.6], we have that $\beta:X\rightarrow X$ is Borel. ∎
For two Borel fields of separable structures $H=(H_{z})_{z\in Z}$ and $K=(K_{z})_{z\in Z}$ and the canonical projections $\pi_{H}:H\rightarrow Z$ and $\pi_{K}:K\rightarrow Z$, we denote by $H\times_{\pi}K$ the Borel set $\{(h,k)\in H\times K\;|\;\pi_{H}(h)=\pi_{K}(k)\}$. The proof of the following corollary relies heavily on the results obtained in [VW25].
@@
We claim that $\phi(\mathfrak{G})$ is a Borel subset of $Y\times X^{\mathbb{N}\times\mathbb{N}}$. To see this, let $\mathcal{S}$ be the set of finite words generated by the elements $\alpha$, $\beta$ and their inverses. By Lemma 3.3, and the fact that composition and inverses of Borel bijections are Borel, we have that if $g\in\mathcal{S}$, then $g:X\rightarrow X$ is Borel. Now consider the Borel map $B_{g}:Y\rightarrow X^{\mathbb{N}\times\mathbb{N}}$ given by $U\mapsto g\cdot(\theta_{i,j}(U))_{i,j}$. For each $g\in\mathcal{S}$, since $B_{g}$ is a Borel map, the graph of $B_{g}$ is Borel (for example see [Tak02, Lemma A.9]). Now since $\mathcal{S}$ is countable, we have that the union of the graphs of $\{B_{g}\;|\;g\in\mathcal{S}\}$ is a Borel subset of $Y\times X^{\mathbb{N}\times\mathbb{N}}$, which is precisely equal to $\phi(\mathfrak{G})$. Now by [VW25, Lemma 4.1], there is a unique standard Borel structure on $\mathfrak{G}$ making the maps $\pi:\mathfrak{G}\rightarrow Y$ and the map $C:\mathfrak{G}\times_{\pi}X\rightarrow X$ Borel. By [VW25, Proposition 4.2], we have that every element of $\mathcal{S}$ is a Borel section $Y\rightarrow\mathfrak{G}$. It follows that the multiplication, inverse and identity maps are Borel and that $\mathfrak{G}=(\mathfrak{G}_{U})_{U\in Y}$ is a Borel field of countable discrete groups. ∎
@@ 4 Construction of genuine ergodic groupoids
Recall the notion of a basis of a discrete measured groupoid from [BCDK24, Definition 3.2]. Roughly speaking, the idea is that a discrete measured groupoid admits a countable disjoint collection of non-null Borel bisections such that the union is the entire groupoid up to measure zero. It turns out that for ergodic groupoids, one can construct a basis consisting of global bisections. This can be seen for example by [Kec95, Exercise 18.15] and we give a proof of this in our context as the next proposition. Notice that for non-ergodic groupoids this cannot be true. For example, if we have a disjoint union of two measured fields of groups consisting of finite and infinite groups almost everywhere respectively, then such a discrete measured groupoid can never admit a basis of global bisections.
@@
Proof.
As in [BCDK24, Proposition 3.1], let $\mathcal{B}=\{X=B_{0},B_{1},B_{2},B_{3},$...\} be a basis of Borel bisections for the groupoid $\mathcal{G}$. We will use ergodicity to ‘patch up’ these bisections and construct global bisections in the process. For convenience, let us label the Borel sets $\mathcal{G}_{E}=\{g\in\mathcal{G}\;|\;s(g)\in E\}$, $\mathcal{G}^{E}=\{g\in\mathcal{G}\;|\;t(g)\in E\}$ and $\mathcal{G}^{F}_{E}=\{g\in\mathcal{G}\;|\;s(g)\in E,t(g)\in F\}$ for Borel subsets $E,F\subseteq X$. We construct a new set of Borel bisections $\mathcal{V}=\{V_{0},V_{1},V_{2},$...\} of $\mathcal{G}$. First let $V_{0}=B_{0}=X$ and suppose that $V_{1},V2,...,V_{i}-1$ are constructed, we shall now construct $V_{i}$. Let $\widetilde{V}=\bigsqcup_{k=1}^{i-1}$ and let $\V~V_{i}^{i}=B_{i}\backslash\widetilde${V}. For $j>i$, recursively define the following Borel sets:
$E^{j}_{i}$ $=X\backslash s(\bigsqcup_{k=i}^{j-1}V^{k}_{i})$
$F^{j}_{i}$ $=X\backslash t(\bigsqcup_{k=i}^{j-1}V^{k}_{i})$
[…]
Notice that by construction and because the maps $s$ and $t$ are countable-to-one, all of the above sets are Borel. Notice also that $V_{i}^{j}$ and $V_{i}^{k}$ are disjoint up to null sets if $j\neq k$. Now we define $V_{i}\coloneqq\bigsqcup_{j\geq i}V_{i}^{j}$ and notice that $V_{i}$ is still a bisection as $V_{i}^{j}$ and $V^{k}_{i}$ have disjoint sources and ranges for all $j\neq k$. Notice that by construction we also have that $V_{i}$’s are mutually disjoint up to null sets, and since $\bigsqcup_{n\geq 0}B_{n}\subseteq\bigsqcup_{n\geq 0}V_{n}$, we even have that $\mathcal{V}$ is a basis for $\mathcal{G}$.
[…]
@@
We first prove that 2 implies 1 by showing that if $\mu$ is atomic, then any discrete measured groupoid with associated equivalence relation $\mathcal{R}$ is a transformation groupoid. Since $\mathcal{R}$ is ergodic and $(X,\mu)$ is atomic, we can assume without loss of generality that $\mathcal{R}$ is the full equivalence relation on an at most countable set $X$ and $\mu$ is the counting measure. Consequently, the isotropy groups $\Gamma_{x}$ are all isomorphic and let us denote this group by $\Gamma$.
Since $\mathcal{G}$ is amenable, by [PSV20, Proposition 6.5], $\mathcal{G}$ can be written as a semi-direct product $\mathcal{R}\ltimes_{\delta}(\Gamma_{x})_{x\in X}$ where $\delta$ is an action of $\mathcal{R}$ on the measured field of groups $(\Gamma_{x})_{x\in X}$. Let $G$ be a group with the same cardinality as $X$ and suppose for convenience the unit space of $\mathcal{G}$ is $G$ with the counting measure. Consider the action of $G\times\Gamma$ on $G$ given by $(h,\gamma)\cdot g=hg$. Consider the transformation groupoid $(G\times\Gamma)\ltimes G$ and the map $\phi:(G\times\Gamma)\ltimes G\rightarrow\mathcal{G}$ given by:
@@
We now check that that $\phi$ is indeed a groupoid homomorphism. First note that:
$(hg,(k,\lambda))(g,(h,\gamma))$ $=(g,(kh,\lambda\gamma))\text{ and }$
$\phi(g,(kh,\lambda\gamma))$ $=((khg,g),\delta_{(g,e)}(\lambda\gamma))$
One can check now that:
$\phi(hg,(k,\lambda))\phi(g,(h,\gamma))$ $=((khg,hg),\delta_{(hg,e)}(\lambda))((hg,g),\delta_{(g,e)}(\gamma))$
@@
It is easy to note now that $\phi$ is a bijective and since the multiplication and the action $\delta$ is Borel, $\phi$ is a Borel groupoid isomorphism.
Now we prove that 1 implies 2, essentially using Proposition 4.1 and Corollary 3.4. Suppose that $\mathcal{R}$ is an ergodic equivalence relation on the diffuse standard probability space $(X,\mu)$ and as in Proposition 4.1 consider a countable family of bijective Borel functions $\theta_{n}:X\rightarrow X$ such that the global bisections $B_{n}=\{(\theta_{n}(x),x)\;|\;x\in X\}$ covers all of $\mathcal{R}$. Let $(G^{0}_{x})_{x\in X}$ be a measured field of finitely generated pairwise non-isomorphic groups (for example as in Corollary 3.4). Let $K$ be an infinite group and note that by [AP10, Proposition 1.3.10] the wreath product $G_{x}=G^{0}_{x}\wr K$ is an icc group for all $x\in X$. By Lemma 2.3, the family of groups $(G_{x})_{x\in X}$ is a measured field of discrete countable groups.
@@
To show that 2 implies 1, one simply notes that if $\mathcal{G}$ is a discrete measured groupoid of type I, then almost every ergodic groupoid in the ergodic decomposition of $\mathcal{G}$ is of type I, and hence the statement follows from Theorem 4.2. To show that 1 implies 2, without loss of generality we can assume that $\mathcal{R}=\mathcal{R}_{\text{ns}}$ as in Notation 2.4. Suppose that $(Z,\eta)$ is a standard probability space and $(\mathcal{R}_{z})_{z\in Z}$ is the ergodic decomposition of $\mathcal{R}$, where $\mathcal{R}_{z}$ is an ergodic equivalence relation on a standard probability space $(Z_{0},\nu_{z})$ for a.e. $z\in Z$. If $(Z,\eta)$ is atomic, then $\mathcal{R}$ is a disjoint union of countably many ergodic equivalence relations. It follows once again from Theorem 4.2 that there exists a genuine groupoid with associated equivalence relation $\mathcal{R}$.
Now suppose that $(Z,\eta)$ is diffuse and let $\nu$ be the integral of the $(\nu_{z})_{z\in Z}$ with respect to $\eta$. We know that $\mathcal{R}$ can be written as an equivalence relation on $(Z_{0}\times Z,\nu)$ given by $(w,z)\sim(w^{\prime},z^{\prime})$ if and only if $z=z^{\prime}$ and $(w,w^{\prime})\in\mathcal{R}_{z}$. Consider a measured field $G=(G_{w})_{w\in Z_{0}}$ of discrete countable finitely generated groups such that they are pairwise non-isomorphic almost everywhere, as in Corollary 3.4. Consider the measured field of groups $H=(H_{(w,z)})_{(w,z)\in Z_{0}\times Z}$ defined as $H_{(w,z)}\coloneqq G_{w}$ for all $z\in Z$ and $w\in Z_{0}$. Since $\mathcal{R}$ is not necessarily ergodic, we do not necessarily have a basis of global bisections as in Proposition 4.1. However we can work around this issue as follows.